# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 5.2.3 Managing Library Dependencies
Managing dependencies in Jupyter notebooks is crucial not only to ensure their reproducibility, but also to ensure their functionality. To address this challenge, researchers have developed various tools. For example, SnifferDog is a dependency management tool that restores notebook execution environments by automatically identifying the required libraries and their compatible versions [180]. This tool leverages a comprehensive API bank to analyze Python code and make notebooks executable and reproducible. Similarly, Davos simplifies dependency management by dynamically installing and updating the correct library versions directly within notebooks, resolving common issues such as version mismatches [35]. Another tool, RELANCER, employs an automatic technique that restores the executability of broken Jupyter notebooks by upgrading deprecated API calls to non-deprecated ones [203].
Additionally, a framework that monitors Linux kernel system calls captures specific versions of dependencies to ensure that the computational environment can be precisely recreated across systems, addressing failures caused by missing or incompatible libraries [185]. These tools collectively aim to reduce the manual effort involved in dependency management, improving the reliability of notebooks in diverse computational settings.
@@ 5.2.4 Performance Analysis
Performance analysis in Jupyter notebooks addresses the efficient execution of notebook code. Werner et al. [190] developed a tool for Jupyter notebooks that measures the execution time and memory usage of individual cells. This tool can identify inefficiencies in runtime or the resource consumption of a notebook to detect performance issues and provide detailed feedback, enabling users to pinpoint resource-intensive areas for refining their code. Extending this idea further, Werner et al. [191] introduced JUmPER, which combines coarse-grained performance monitoring with fine-grained instrumentation for interactive HPC workflows. In addition, JUmPER leverages parallel marshalling and in-memory communication methods to minimize instrumentation overhead.
Summary of managing computational environment and workflow Managing the computational environment and workflow in Jupyter notebooks is critical to ensuring smooth execution. Empirical studies show that notebook workflows typically involve iterative processes, from importing data and editing code to generating output and sharing results [78, 124]. Tools, such as directed acyclic graphs and unique cell identifiers, have been introduced to help map workflows and manage dependencies between cells [130]. Researchers developed an efficient computational notebook engine for enabling cell-wise checkpointing [141]. Researchers have also explored solutions for adapting notebooks to different computational environments, including tools for migrating cells to distributed platforms such as Kubernetes [31] and systems such as Elastic-Notebook [88] for checkpointing and restoration. In addition, managing library dependencies remains a key focus, with tools such as RELANCER [203], SnifferDog [180], and Davos [35] automating the resolution of API issues and ensuring consistency across environments. These advancements address the unique challenges of Jupyter notebooks, making them more reliable and scalable.
@@ 5.3 Readability of Notebooks
The readability of notebooks has two central dimensions: (1) the readability of the code in the notebook, which focuses on the clarity, organization and quality of the written code [90], and (2) the readability of the notebook narrative, which emphasizes how effectively the text and the output explain the workflow and analysis [133]. Unlike traditional source code files, notebooks integrate iterative code execution, visual output, and narrative explanations, contributing to a distinctively different in nature narrative-driven coding environment [135]. Researchers have addressed notebook readability challenges through various practices and tools, with the aim of improving both code and narrative readability.
@@ 5.3.1 Refactoring
Refactoring helps enhance readability by focusing primarily on improving the clarity and organization of the notebook’s code itself. Due to Jupyter Notebook’s distinctive characteristics, such as the cell-based structure and the iterative nature of cell-wise execution [90], traditional software refactoring methods may not always be directly applied. Often, notebooks are developed with a focus on completing the final data analysis, and little attention is paid to code refactoring, as users prioritize drawing conclusions and sharing results [135]. This approach leads to “messy” notebooks, which are disorganized and poorly structured, and characterized by scattered cell arrangements, excessive inline outputs, and a lack of modularity [55]. Messy code reduces readability, makes notebooks harder to maintain, and complicates tracing the steps that produced the results, which can ultimately discourage sharing and collaboration [133, 55].
@@
To address these issues, refactoring practices are employed to improve notebook readability and maintainability. Researchers found that the common refactoring practices in notebooks are extracting functions, reordering cells, renaming notebooks, splitting cells, and merging cells [90]. In another study, Dong et al. [30] expanded this list of refactorings by identifying additional activities such as removing commented blocks, adding comments, and renaming variables. These refactoring practices are not universally applied but vary depending on the authors’ backgrounds (e.g., data scientists versus computer scientists) [90]. These also vary from the intended use of the notebook [29]. Dong [29] noticed that notebooks for sharing results with others saw mostly refactorings related to Markdown, whereas, for production notebooks, the most refactorings occurred in the logical code in the transition between notebook and Python file.
@@
In addition, tools such as nbslicer, a hybrid static-dynamic slicing tool, were developed to help cleaning and re-executing notebooks more effectively [146]. Using backward and forward program slicing, nbslicer helps reduce messy code and improves the efficiency of refactoring while addressing challenges such as achieving accurate program slices without excessive overhead or performance issues. These tools and practices collectively aim to improve the readability and maintainability of Jupyter notebooks. Another tool named ReSplit helps to split the notebook code cell so that coherent code lines are placed together [159]. In the first step, the algorithm of that tool suggests merging some of the cells, and in the second step, it suggests extracting specific code fragments from the original cells into new ones [159].
@@
Refactoring should not affect the functionality of the code. However, verifying the correctness of a notebook after refactoring can be challenging. For example, when code clones are refactored into functions, users must ensure that the output remains consistent. To address this issue, [140] proposed a method to verify correctness by comparing API calls and textual output to ensure that the same API calls with identical parameters produce consistent results, simplifying the validation process.
@@ 5.3.2 Nonlinear Visualization of Notebooks
While refactoring primarily focuses on code readability, nonlinear visualization focuses on improving readability through narrative clarity and interactive structuring. Due to the inherent cell-based organization of Jupyter notebooks which follow a linear, top-to-bottom format, the narrative of notebooks can sometimes limit readability. Researchers addressed narrative readability by proposing nonlinear organization tools that enable visualization and organization beyond the linear convention. For example, tools such as StickyLand [183] and ToonNote [64] allow users to rearrange notebook cells based on their priorities, such as by grouping related sections or highlighting key parts of the workflow. These tools offer new options for organizing, navigating, and displaying notebooks.
StickyLand [183] provides a visual interface where users can freely drag and drop cells to create a customized layout, breaking away from the conventional top-to-bottom order. This allows users to better align the notebook’s structure with their thought process or the natural flow of their analysis. Figure 6 demonstrates how StickyLand transforms the notebook interface into a more intuitive and visually organized workspace.
@@
Furthermore, Chattopadhyay et al. [20] explored the cognitive processes involved in understanding computational notebooks, identifying key tasks such as comprehension, mental modelling, and contextual inference. They mapped these tasks to practical design elements, including navigation panels, annotations, and structured titles, and incorporated these elements into Porpoise, an interactive overlay tool specifically designed to enhance notebook sense making and navigation.
@@
Summary of readability of notebooks Notebook readability refers to readability and organization of code and other content within Jupyter notebooks. Researchers have examined refactoring practices to improve code readability, such as reorganizing cells, extracting functions, and splitting or merging cells. Tools such as nbslicer [146] and ReSplit [159] support these efforts by providing semi-automated solutions to clean up and optimize notebook structures. Moreover, tools such as StickyLand [183] and ToonNote [64] facilitate non-linear visualization of notebooks, enabling users to organize cells and better reflect their analytical processes dynamically. Collectively, these advancements improve the readability and maintainability of Jupyter notebooks.
@@ 5.4 Documentation of Notebooks
Research on notebook documentation is crucial because Jupyter Notebook allows combining code with narrative, data visualization, and exploratory analysis in a way that traditional source code files do not [72]. However, notebook users often pay attention only to the code without creating or updating their documentation [174]. For our purposes, we define documentation to include any content within Markdown cells and comments in code cells. In this section, we will discuss the key areas that researchers have explored regarding notebook documentation, including document generation, cell header creation, and storytelling.
@@ 5.4.1 Empirical Studies on Documentation
Integrating explanatory text in notebook documentation is essential to improve the understandability and shareability of notebooks. These texts make a well-documented notebook that describes workflows, provides context, and explains the purpose of code. Although studies have shown that almost all notebooks (99%) include at least one Markdown cell [135], the quality of documentation inside markdowns is not up to mark [172]. To improve notebook documentation, Wang et al. [172] analyzed highly voted notebooks and identified nine guidelines for Markdown content (such as adding sections or subsections) in notebooks. Researchers also showed that although using code comments can enhance readability by organizing and annotating the flow of code [125], only a small fraction of notebooks use comments to explain the reasoning (10%) or expected results (4%) of the corresponding code cells [135, 136].
@@ 5.4.2 Documentation Generation
Researchers developed several tools to generate notebook documentation. To tackle the challenge of insufficient documentation in data science notebooks, Themisto, an AI system to generate code documentation, has been proposed [174, 169]. It works in three phases, i.e., retrieving relevant API documentation, generating documentation automatically, and prompting users to add their documentation. Themisto was reported to reduce documentation time, increase user satisfaction, and encourage documentation of previously overlooked code. Another tool HAConvGNN, uses a hierarchical attention mechanism to focus on relevant code snippets and tokens for accurate documentation generation [91]. In addition, Cell2Doc uses ML pipelines to generate documentation for code cells [100] and InkSight [89] documents insights from charts using sketch-based interactions. To enhance the quality and relevance of automatically generated documentation, Ghahfarokhi et al. [41] proposed a CNN-RNN-based approach leveraging diverse code metrics (such as cell complexity and API popularity). Empirical results from Muller et al. [101] show that in 41% of the cases, users still need to modify automatically generated notebook documentation.
Storytelling for Jupyter notebooks takes documentation even further by transforming raw analyses into coherent narratives tailored for communication. For instance, Notable [81] automates the generation of narrative presentation slides directly from the notebooks. Addressing the related challenge of slide creation from disorganized computational notebooks, Wang et al. [175] introduced OutlineSpark to generate slides through interactive outlining and computational support. Going one step further, Ouyang et al. [107] proposed NotePlayer, which bridges notebook cells with dynamic video segments by integrating computational engines and LLM-generated narrations, addressing the problem of excessive manual effort traditionally required for detailed tutorial video creation.
@@ 5.4.3 Cell Header Generation
In notebooks, a cell header is typically a Markdown cell that defines section titles or headings, which helps organize and structure the notebook more effectively. Cell headers often utilize Markdown syntax (e.g., # or ##) to create hierarchical headers, enhancing the readability and navigability. Unlike standard documentation generation, which focuses on providing detailed explanations, inputs, or outputs, cell headers primarily serve to visually structure the flow of the notebook [165]. These headers act as navigational aids, dividing the notebook into logical sections to improve readability and guide users through the workflow. Venkatesh et al. [167, 166] introduced an automatic cell header generation tool HeaderGen, which analyzes the functions and call graphs of the notebook code to generate appropriate headers for the notebook cells.  Venkatesh and Bodden [165] published an early prototype of the HeaderGen tool that can automatically create headers in Markdown cells using static analysis of the code cells, even without providing specific details about the Markdown text documentation. Extending beyond static analysis, Perez et al. [111] proposed a hybrid approach called JUPYLABEL, which incorporated both rule-based heuristics and decision-tree classifiers to classify and label notebook cells to generate headers.
Summary of documentation of notebooks Research on notebook documentation in Jupyter notebooks addressed integrating explanatory text to enhance understandability and shareability. In documentation generation, ML models help automate the process, with tools such as Themisto [174] and HAConvGNN [91] which use ML to generate comprehensive documentation based on code structure and content. Research has explored human interaction with AI-generated text, highlighting the need for human intervention to improve the quality of generated documentation. Cell header generation focuses on automatically generating headers to improve notebook navigability and usability, exemplified by tools such as HeaderGen [167] and JUPYLABEL [111].

### v2

@@ 5.2.3 Managing Library Dependencies
Managing dependencies in Jupyter notebooks is crucial not only to ensure their reproducibility, but also to ensure their functionality. To address this challenge, researchers have developed various tools. For example, SnifferDog is a dependency management tool that restores notebook execution environments by automatically identifying the required libraries and their compatible versions [wang2021restoring]. This tool leverages a comprehensive API bank to analyze Python code and make notebooks executable and reproducible. Similarly, Davos simplifies dependency management by dynamically installing and updating the correct library versions directly within notebooks, resolving common issues such as version mismatches [fitzpatrick2022davos]. Another tool, RELANCER, employs an automatic technique that restores the executability of broken Jupyter notebooks by upgrading deprecated API calls to non-deprecated ones [zhu2021restoring].
Additionally, a framework that monitors Linux kernel system calls captures specific versions of dependencies to ensure that the computational environment can be precisely recreated across systems, addressing failures caused by missing or incompatible libraries [wannipurage2022framework]. These tools collectively aim to reduce the manual effort involved in dependency management, improving the reliability of notebooks in diverse computational settings.
@@ 5.2.4 Performance Analysis
Performance analysis in Jupyter notebooks addresses the efficient execution of notebook code. werner2021bridging developed a tool for Jupyter notebooks that measures the execution time and memory usage of individual cells. This tool can identify inefficiencies in runtime or the resource consumption of a notebook to detect performance issues and provide detailed feedback, enabling users to pinpoint resource-intensive areas for refining their code. Extending this idea further, werner2024jumper introduced JUmPER, which combines coarse-grained performance monitoring with fine-grained instrumentation for interactive HPC workflows. In addition, JUmPER leverages parallel marshalling and in-memory communication methods to minimize instrumentation profiling overhead even in distributed computing scenarios. Beyond instrumentation and profiling, other research has explored predictive and benchmarking approaches. For example, prathanrat2018performance utilized machine learning models to predict the performance of notebooks running on multi-user JupyterHub servers, helping administrators anticipate potential resource bottlenecks. faenza2024containerized examined the impact of containerizing Jupyter notebooks on execution performance, evaluating how the isolation and flexibility provided by containers trade off against any added overhead. For educational use cases, oden2024integrating developed an approach to integrate interactive performance analysis into notebooks used in parallel programming courses, allowing students to visualize how code changes affect execution time and resource usage in real time. Additionally, grotov2025themisto proposed Themisto, a Jupyter-based benchmarking framework to systematically measure and compare the runtime performance of notebook code under different conditions. These contributions complement cell-level profilers by offering higher-level insights into notebook efficiency, thereby guiding users in optimizing their notebooks for better performance.
Summary of managing computational environment and workflow Managing the computational environment and workflow in Jupyter notebooks is critical to ensuring smooth execution. Empirical studies show that notebook workflows typically involve iterative processes, from importing data and editing code to generating output and sharing results [lau2020design, raghunandan2023measuring]. Tools, such as directed acyclic graphs and unique cell identifiers, have been introduced to help map workflows and manage dependencies between cells [rehman2019towards]. Researchers developed an efficient computational notebook engine for enabling cell-wise checkpointing [sato2024multiverse]. Researchers have also explored solutions for adapting notebooks to different computational environments, including tools for migrating cells to distributed platforms such as Kubernetes [duan2023jup2kub] and systems such as Elastic-Notebook [li2023elasticnotebook] for checkpointing and restoration. In addition, managing library dependencies remains a key focus, with tools such as RELANCER [zhu2021restoring], SnifferDog [wang2021restoring], and Davos [fitzpatrick2022davos] automating the resolution of API issues and ensuring consistency across environments. These advancements address the unique challenges of Jupyter notebooks, making them more reliable and scalable.
@@ 5.3 Readability of Notebooks
The readability of notebooks has two central dimensions: (1) the readability of the code in the notebook, which focuses on the clarity, organization and quality of the written code [liu2023refactoring], and (2) the readability of the notebook narrative, which emphasizes how effectively the text and the output explain the workflow and analysis [rule2018ten]. Unlike traditional source code files, notebooks integrate iterative code execution, visual output, and narrative explanations, contributing to a distinctively different in nature narrative-driven coding environment [rule2018exploration]. Researchers have addressed notebook readability challenges through various practices and tools, with the aim of improving both code and narrative readability.
@@ 5.3.1 Refactoring
Refactoring helps enhance readability by focusing primarily on improving the clarity and organization of the notebook’s code itself. Due to Jupyter Notebook’s distinctive characteristics, such as the cell-based structure and the iterative nature of cell-wise execution [liu2023refactoring], traditional software refactoring methods may not always be directly applied. Often, notebooks are developed with a focus on completing the final data analysis, and little attention is paid to code refactoring, as users prioritize drawing conclusions and sharing results [rule2018exploration]. This approach leads to “messy” notebooks, which are disorganized and poorly structured, and characterized by scattered cell arrangements, excessive inline outputs, and a lack of modularity [head2019managing]. Messy code reduces readability, makes notebooks harder to maintain, and complicates tracing the steps that produced the results, which can ultimately discourage sharing and collaboration [rule2018ten, head2019managing].
@@
To address these issues, refactoring practices are employed to improve notebook readability and maintainability. Researchers found that the common refactoring practices in notebooks are extracting functions, reordering cells, renaming notebooks, splitting cells, and merging cells [liu2023refactoring]. In another study, dong2021splitting expanded this list of refactorings by identifying additional activities such as removing commented blocks, adding comments, and renaming variables. These refactoring practices are not universally applied but vary depending on the authors’ backgrounds (e.g., data scientists versus computer scientists) [liu2023refactoring]. These also vary from the intended use of the notebook [dong2021qualitative]. dong2021qualitative noticed that notebooks for sharing results with others saw mostly refactorings related to Markdown, whereas, for production notebooks, the most refactorings occurred in the logical code in the transition between notebook and Python file.
@@
In addition, tools such as nbslicer, a hybrid static-dynamic slicing tool, were developed to help cleaning and re-executing notebooks more effectively [shankar2022bolt]. Using backward and forward program slicing, nbslicer helps reduce messy code and improves the efficiency of refactoring while addressing challenges such as achieving accurate program slices without excessive overhead or performance issues. These tools and practices collectively aim to improve the readability and maintainability of Jupyter notebooks. Another tool named ReSplit helps to split the notebook code cell so that coherent code lines are placed together [titov2022resplit]. In the first step, the algorithm of that tool suggests merging some of the cells, and in the second step, it suggests extracting specific code fragments from the original cells into new ones [titov2022resplit].
@@
Refactoring should not affect the functionality of the code. However, verifying the correctness of a notebook after refactoring can be challenging. For example, when code clones are refactored into functions, users must ensure that the output remains consistent. To address this issue, [sato2022comparing] proposed a verification method by comparing API calls and textual output to ensure that the same API calls with identical parameters produce consistent results.
@@ 5.3.2 Nonlinear Visualization of Notebooks
While refactoring primarily focuses on code readability, nonlinear visualization focuses on improving readability through narrative clarity and interactive structuring. Due to the inherent cell-based organization of Jupyter notebooks which follow a linear, top-to-bottom format, the narrative of notebooks can sometimes limit readability. Researchers addressed narrative readability by proposing nonlinear organization tools that enable visualization and organization beyond the linear convention. For example, tools such as StickyLand [wang2022stickyland] and ToonNote [kang2021toonnote] allow users to rearrange notebook cells based on their priorities, such as by grouping related sections or highlighting key parts of the workflow. These tools offer new options for organizing, navigating, and displaying notebooks. StickyLand [wang2022stickyland] provides a visual interface where users can freely drag and drop cells to create a customized layout, breaking away from the conventional top-to-bottom order. This allows users to better align the notebook’s structure with their thought process or the natural flow of their analysis. Figure 6 demonstrates how StickyLand transforms the notebook interface into a more intuitive and visually organized workspace.
Building on the idea of enhancing notebook interactivity and structure, grotov2025evolving presented the notebook interface as a two-dimensional space, allowing users to spatially organize code cells, detach outputs, and create separate environments for experimentation. Other researchers identified three main 2D cell arrangement patterns in computational notebooks, namely Linear, Multi-Column, and Workboard, highlighting their value for branching and comparative analyses [harden2022exploring, harden2023there]. Additionally, rawn2025pagebreaks proposed Pagebreaks, a language construct that introduces scoped regions across multiple notebook cells. Unlike traditional functions, Pagebreaks preserve exploratory behaviors such as interleaving code and output or composing programs incrementally, while mitigating confusion caused by global variables. husak2023slicito developed a tool to help developers by interactively explore project hierarchies and slice relevant code segments, offering a moldable and task-specific view of complex systems.
@@
Furthermore, chattopadhyay2023make explored the cognitive processes involved in understanding computational notebooks, identifying key tasks such as comprehension, mental modeling, and contextual inference. They mapped these tasks to practical design elements, including navigation panels, annotations, and structured titles, and incorporated these elements into Porpoise, an interactive overlay tool specifically designed to enhance notebook sense making and navigation. Additionally, merino2022making introduced interactive widgets built on an exploring interpreter that visualize execution history and variable states. These widgets allow users to access previous execution across configurations and reduce cognitive load by making program state transitions explicit.
5.3.3 Accessibility
Computational notebooks often lack structural features, such as headings, alternative text, and accessible visual outputs, that are essential for blind and visually impaired users to navigate, comprehend, and interact with content via screen readers. To address this issue, venkatesh2023notably conducted a large-scale analysis of 100,000 Python-based Jupyter notebooks to identify systemic accessibility barriers in authoring practices, data representation, and publishing workflows. Their findings revealed that most notebooks are structurally inaccessible: 57% lacked output cells, 43% contained figures without descriptive alternatives, and semantic elements like headings and tables were underutilized. Additionally, charts and other visual outputs were rarely accompanied by accessible counterparts, such as explanatory text or tabular summaries. To improve accessibility, nylund2025matplotalt introduced a Python library called MatplotAlt, which enables users to add alternative text (alt text) to Matplotlib figures in computational notebooks. Building on this, we found an urgent need for accessibility aware design in notebook documentation to ensure inclusive practices for notebook developers, including adding alternative text and generating accessible tables alongside visualizations to support screen reader navigation.
@@ 5.3.3 Accessibility
Summary of readability of notebooks Notebook readability refers to readability and organization of code and other content within Jupyter notebooks. Researchers have examined refactoring practices to improve code readability, such as reorganizing cells, extracting functions, and splitting or merging cells. Tools such as nbslicer [shankar2022bolt] and ReSplit [titov2022resplit] support these efforts by providing semi-automated solutions to clean up and optimize notebook structures. Moreover, tools such as StickyLand [wang2022stickyland], ToonNote [kang2021toonnote], Pagebreaks [rawn2025pagebreaks] facilitate non-linear visualization of notebooks, enabling users to organize cells and better reflect their analytical processes dynamically. Additionally, accessibility remains a critical concern in notebook documentation, with widespread structural barriers limiting screen reader usability and highlighting the need for inclusive authoring practices [venkatesh2023notably, nylund2025matplotalt]. Collectively, these advancements improve the readability and maintainability of Jupyter notebooks.
@@ 5.4 Documentation of Notebooks
Research on notebook documentation is crucial because Jupyter Notebook allows combining code with narrative, data visualization, and exploratory analysis in a way that traditional source code files do not [kluyver2016jupyter]. However, notebook users often pay attention only to the code without creating or updating their documentation [wang2022documentation]. For our purposes, we define documentation to include any content within Markdown cells and comments in code cells. In this section, we will discuss the key areas that researchers have explored regarding notebook documentation, including document generation, cell header creation, and storytelling.
@@ 5.4.1 Empirical Studies on Documentation
Integrating explanatory text in notebook documentation is essential to improve the understandability and shareability of notebooks. These texts make a well-documented notebook that describes workflows, provides context, and explains the purpose of code. Although studies have shown that almost all notebooks (99%) include at least one Markdown cell [rule2018exploration], the quality of documentation inside markdowns is not up to mark wang2021makes. To improve notebook documentation, wang2021makes analyzed highly voted notebooks and identified nine guidelines for Markdown content (such as adding sections or subsections) in notebooks. Researchers also showed that although using code comments can enhance readability by organizing and annotating the flow of code [raghunandan2022code], only a small fraction of notebooks use comments to explain the reasoning (10%) or expected results (4%) of the corresponding code cells [rule2018exploration, rule2018design]. Complementing studies on Markdown and code comments, wong2025method found that method names in Jupyter notebooks often lack clarity as embedded documentation, with only 55.57% starting with a verb and 30.39% containing abbreviations or acronyms. These naming practices reduce the effectiveness of method names as notebook documentation, underscoring the need for clearer naming conventions in notebook environments.
@@ 5.4.2 Documentation Generation
Researchers developed several tools to generate notebook documentation. To tackle the challenge of insufficient documentation in data science notebooks, Themisto, an AI system to generate code documentation, has been proposed [wang2022documentation, wang2021graph]. It works in three phases, i.e., retrieving relevant API documentation, generating documentation automatically, and prompting users to add their documentation. Themisto was reported to reduce documentation time, increase user satisfaction, and encourage documentation of previously overlooked code. Another tool HAConvGNN, uses a hierarchical attention mechanism to focus on relevant code snippets and tokens for accurate documentation generation [liu2021haconvgnn]. In addition, schmieg2025enhancing showed that well-structured notebooks with detailed explanations, dataset descriptions, and visual support enhance knowledge transfer and onboarding in machine learning research. Cell2Doc uses ML pipelines to generate documentation for code cells [mondal2023cell2doc] and InkSight [lin2023inksight] documents insights from charts using sketch-based interactions.
To enhance the quality and relevance of automatically generated documentation, a CNN-RNN-based approach has been proposed that leverages diverse code metrics such as cell complexity and API popularity [mostafavi2024beyond]. In addition, incorporating code metrics such as complexity and API usage into deep learning models significantly improves the accuracy of auto-generated documentation, with BLEU and ROUGE scores showing consistent gains [mostafavi2024can]. Complementing these efforts, a plugin called InterLink was introduced to visually link text, code, and outputs in a side-by-side layout, thereby improving the traceability of notebook documentation [lin2025interlink]. Empirical results from muller2021data show that in 41% of the cases, users still need to modify automatically generated notebook documentation.
Storytelling for Jupyter notebooks takes documentation even further by transforming raw analyses into coherent narratives tailored for communication. For instance, Notable [li2023notable] automates the generation of narrative presentation slides directly from the notebooks. Addressing the related challenge of slide creation from disorganized computational notebooks, wang2024outlinespark introduced OutlineSpark to generate slides through interactive outlining and computational support. Going one step further, ouyang2024noteplayer proposed NotePlayer, which bridges notebook cells with dynamic video segments by integrating computational engines and LLM-generated narrations, addressing the problem of excessive manual effort traditionally required for detailed tutorial video creation.
@@ 5.4.3 Cell Header Generation
In notebooks, a cell header is typically a Markdown cell that defines section titles or headings, which helps organize and structure the notebook more effectively. Cell headers often utilize Markdown syntax (e.g., # or ##) to create hierarchical headers, enhancing the readability and navigability. Unlike standard documentation generation, which focuses on providing detailed explanations, inputs, or outputs, cell headers primarily serve to visually structure the flow of the notebook [venkatesh2021automated]. These headers act as navigational aids, dividing the notebook into logical sections to improve readability and guide users through the workflow. venkatesh2023enhancing, venkatesh2024static introduced an automatic cell header generation tool HeaderGen, which analyzes the functions and call graphs of the notebook code to generate appropriate headers for the notebook cells.  venkatesh2021automated published an early prototype of the HeaderGen tool that can automatically create headers in Markdown cells using static analysis of the code cells, even without providing specific details about the Markdown text documentation. Extending beyond static analysis, perez2024flexible proposed a hybrid approach called JUPYLABEL, which incorporated both rule-based heuristics and decision-tree classifiers to classify and label notebook cells to generate headers.
Summary of documentation of notebooks Research on notebook documentation in Jupyter notebooks addressed integrating explanatory text to enhance understandability and shareability. In documentation generation, ML models help automate the process, with tools such as Themisto [wang2022documentation] and HAConvGNN [liu2021haconvgnn] which use ML to generate comprehensive documentation based on code structure and content. Research has explored human interaction with AI-generated text, highlighting the need for human intervention to improve the quality of generated documentation. Cell header generation focuses on automatically generating headers to improve notebook navigability and usability, exemplified by tools such as HeaderGen [venkatesh2023enhancing] and JUPYLABEL [perez2024flexible].